OpenStudy (anonymous):

Find the 6*1/4% of Rs.2400.

OpenStudy (anonymous):

6/4 *1/100 * 2400 = 1

OpenStudy (anonymous):

can u explain with the diagram.

OpenStudy (anonymous):

no its 6*1/4 = 25/4

OpenStudy (anonymous):

its a mixed fraction

ganeshie8 (ganeshie8):

\(6\frac{1}{4} \% \ of\ 2400 = ? \)

OpenStudy (anonymous):

oh...i misread the question

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

25/4 %

OpenStudy (anonymous):

change the mixed fraction

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

the same way which i used first

OpenStudy (anonymous):

25/4 / 100 * 2400

OpenStudy (anonymous):